### v1

# Primes and polygonal numbers

Abstract.
For all integers $r,s>2$, there are infinitely many primes which can be written as the sum of an $r$-gonal number and an $s$-gonal number. In general, any linear combination of an $r$-gonal number and an $s$-gonal number with coprime positive integer coefficients produces infinitely many primes. However, unless $r=s=4$, such primes can not be described by finitely many congruences.
Key words and phrases:
prime, polygonal number, triangular number, arithmetic progression
2020 Mathematics Subject Classification:
Primary 11N32, 11N13, 11N36; Secondary 11P32, 11B25, 11B05
1. Introduction and the main result
A prime is a sum of two squares if and only if it is of the form $14k+$ 1 (see [8]). Since there are infinitely many primes of the form $14k+$ 1, it follows that there are infinitely many primes which are sum of two squares. It is natural to wonder whether the same holds for triangular numbers or more generally, whether for all integers $r,s\geq 3$, there are infinitely many primes which can be represented as the sum of an $r$-gonal number and an $s$-gonal number, where the $m$-th $r$-gonal number is defined by (see Chapter 2 in [2])
$T(r,m):=(r-2)\frac{m(m-1)}{2}+m.$ (1.1)
Even more generally, one may ask whether for each pair of integers $r,s\geq 3$ and for each pair of mutually coprime positive integers $\mu$ and $\nu$, there are infinitely many primes of the form
$\mu T(r,m)+\nu T(s,n).$
We answer this in affirmative.
Theorem 1.
For sufficiently large $N$, for each pair of integers $r,s\geq 3$ and for each pair of mutually coprime positive integers $\mu$ and $\nu$, the number of primes of the form
$\mu T(r,m)+\nu T(s,n)$
which are less than $N$ is
$\asymp\begin{cases}\dfrac{N}{\log N},&\text{ if }r=s=4\\ \\ \dfrac{N}{(\log N)^{3/2}}&\text{otherwise.}\end{cases}$
By an admissible arithmetic progression, we mean an integer arithmetic progression that has a pair of mutually coprime terms.
Corollary 1.
Unless $r=s=4$, for each pair of mutually coprime positive integers $\mu$ and $\nu$, the primes of the form
$\mu T(r,m)+\nu T(s,n),$
can not be described as the primes contained in a set of admissible arithmetic progressions.
In other words, one can only describe the primes of the form
$ax^{2}+by^{2}$
by a finite set of congruences for certain $a,b\in\mathbb{N}$ (see [3]). However, Theorem 1 together with the Prime Number Theorem for arithmetic progressions (Corollary 11.20 in [7]) imply that for any pair of $a,b\in\mathbb{N}$, if we replace either $x^{2}$ or $y^{2}$ above by another polygonal number, the resulting linear combination of polygonal numbers misses infinitely many primes in every admissible arithmetic progression.
In particular, the sequence $\{4k+1\}_{k\in\mathbb{N}}$ or its subprogressions are the only admissible arithmetic progressions such that each prime in those progressions is a sum of two polygonal numbers.
It follows from Theorem 1 and Chebyshev’s theorem (Corollary 2.6 in [7]) that the set of the primes, which are sum of two polygonal numbers (or more generally, linear combination of two polygonal numbers with coprime positive integer coefficients) at least one of which is not a square, does not have positive relative upper density with respect to the set of all primes. So, Green-Tao Theorem [4] which guarantees the existence of arbitrarily long arithmetic progressions among the subsets of the primes with positive relative upper densities, is not applicable here either.
From Theorem 1, for $\mu=\nu=1$ and $r=s=3$, we obtain the following corollary which answers the Open Problem 1 in [1].
Corollary 2.
There are infinitely many primes which are sum of two triangular numbers.
Since a number $n$ is a sum of two triangular numbers if and only if $14n+$ 1 is a sum of two squares (see [1]), we have
Corollary 3.
There are infinitely many primes $p$ such that $14p+$ 1 is a sum of two squares.
It is natural to wonder whether there are infinitely many primes $p$ such that both $p$ and $14p+$ 1 are sums of two squares or more generally, whether both the residue classes $1$ and $3$ modulo $4$ contain infinitely many primes $p$ such that $14p+$ 1 is a sum of two squares. Even more generally, one may ask:
Open Question 1.
Does every admissible arithmatic progression contain infinitely many primes which are sum of two triangular numbers?
2. Proof of the main result
The main ingredients of the proof of Theorem 1 are a couple of results by Iwaniec [5], which we state below.
Theorem A (Theorem 1 in [5]).
A Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g$ be of an irreducible polynomial of degree $2$ in $\mathbb{Z}[x,y]$ such that $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly independent and $P(x,y)$ represents arbitrarily large odd numbers for $x,y\in\mathbb{N}$. Then
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\gg\frac{N}{\log N},$ (2.1)
if $\Delta:=b^{2}-4ac$ is a perfect square or $D:=af^{2}-bef+ce^{2}+g\Delta=0$, and
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\asymp\frac{N}{(\log N)^{3/2}},$ (2.2)
if $\Delta$ is not a perfect square and $D\neq 0$.
Lemma B (Lemma 1 in [5]).
B Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g\in\mathbb{Z}[x,y]$. Then $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly dependent if and only if
$\Delta=b^{2}-4ac=0,\ \alpha:=bf-2ce=0\text{ \ and \ }\beta:=be-2af=0.$
Proof of Theorem 1: For a pair of integers $r,s\geq 3$ and for a pair of mutually coprime integers $\mu,\nu>0$, considering without loss of generality all possible parities of the natural numbers $m$ and $n$ in
$\mu T(r,m)+\nu T(s,n),$
we note that (see (1.1)) the above linear combination is equal to
$P_{1}(x,y):=2\mu(r-2)x^{2}+2\nu(s-2)y^{2}+\mu(4-r)x+\nu(4-s)y$ (2.3)
for $m=2x$ and $n=2y$,
$P_{2}(x,y):=2\mu(r-2)x^{2}+2\nu(s-2)y^{2}+\mu(4-r)x+\nu sy+\nu$ (2.4)
for $m=2x$ and $n=2y+1$,
$P_{3}(x,y):=2\mu(r-2)x^{2}+2\nu(s-2)y^{2}+\mu rx+\nu sy+\mu+\nu$ (2.5)
for $m=2x+1$ and $n=2y+1$.
Since the coefficients of $x^{2}$ and $y^{2}$ in each of the above polynomials are positive and since the coefficient of $xy$ is zero, it follows that
$\Delta<0$ (2.6)
for each of the above polynomials (see Lemma B). Hence, $\frac{\partial{P_{i}}}{\partial x}$ and $\frac{\partial{P_{i}}}{\partial y}$ linearly independent for all $i\in\{1,2,3\}$.
Also, from (2.3), (2.4) and (2.5), we note that for each of the polynomials $P_{1},P_{2}$ and $P_{3}$, we have
$D=2\mu\nu^{2}(r-2)(4-s)^{2}+2\mu^{2}\nu(s-2)(4-r)^{2},$ (2.7)
which equals zero if and only if $r=s=4$ (see Theorem A).
In order to apply Theorem A, we must also check whether the polynomials $P_{1},P_{2}$ $P_{3}$ are irreducible in $\mathbb{Z}[x,y]$. Suppose, for some $i\in\{1,2,3\}$, there exists $a,b,c,a^{\prime},b^{\prime},c^{\prime}\in\mathbb{Z}$ such that
$P_{i}(x,y)=(ax+by+c)(a^{\prime}x+b^{\prime}y+c^{\prime}).$
Comparing the coefficients of $x^{2}$, $y^{2}$ and $xy$ in both of the above expressions for $P_{i}(x,y)$, we obtain
$aa^{\prime}=2\mu(r-2),\ bb^{\prime}=2\nu(s-2)\text{ and }ab^{\prime}+a^{\prime}b=0.$
Since $\mu,\nu>0$ and $r,s>2$, the first two equalities imply that $aa^{\prime}>0$ and $bb^{\prime}>0$ and the last equality implies that there exists $k\in\mathbb{Q}$ such that $\frac{a}{b}=-\frac{a^{\prime}}{b^{\prime}}=k$. Hence, substituting $kb$ for $a$ and $-kb^{\prime}$ for $a^{\prime}$ in the first equality and dividing both sides of it with the corresponding sides of the second equality above, we obtain
$k^{2}=-\frac{\mu(r-2)}{\nu(s-2)}<0,$
which is absurd! Thus, $P_{i}(x,y)$ is irreducible in $\mathbb{Q}[x,y]$ for all $i\in\{1,2,3\}$. Hence, it follows from Gauss’s Lemma (see Ch.VI, §2 in [6]) that $P_{1},P_{2}$ and $P_{3}$ are irreducible in $\mathbb{Z}[x,y]$ if and only if these are primitive, i.e. the greatest common divisor of the coefficients of each of these polynomials is 1.
Since for all $i\in\{1,2,3\}$, the coefficients of $x^{2}$ and $y^{2}$ in $P_{i}(x,y)$ are positive and even, whereas the coefficient of $xy$ is zero, it follows that if any of the remaining coefficients of $P_{i}$ are odd, then $P_{i}(x,y)$ produces infinitely many odd numbers for $x,y\in\mathbb{N}$. In particular, if the greatest common divisors of the coefficients of $P_{1}$, $P_{2}$, $P_{3}$ are 1, then each of these polynomials represents arbitrarily large odd numbers. In the following we compute the gcd $G$ of the coefficients of $P_{i}$ for $i\in\{1,2,3\}$.
Case I. For $P_{1}$, from (2.3) we note that
$G=\gcd(2\mu(r-2),2\nu(s-2),\mu(4-r),\nu(4-s))=\gcd(4,\mu r,\nu s).$
Hence, $G=1$ if and only if either $\mu r$ or $\nu s$ is odd. In particular, $G=1$ implies $D>0$ (see (2.7)). So, it follows from Theorem A that
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{1}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\hskip 5.0pt\begin{cases}\asymp\dfrac{N}{(\log N)^{3/2}}&\text{if $\{\mu r,\nu s\}\nsubseteq 2\mathbb{Z}$}\\ \\ \leq 1&\text{otherwise. }\end{cases}$
Case II. For $P_{2}$, from (2.4) we note that
$G=\gcd(2\mu(r-2),2\nu(s-2),\mu(4-r),\nu s,\nu)=\gcd(4,r,\nu).$
Hence, $G=1$ if and only if either $r$ or $\nu$ is odd. So, it follows from (2.7), Theorem A and Chebyshev’s theorem (Corollary 2.6 in [7]) that
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{2}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\hskip 5.0pt\begin{cases}\asymp\dfrac{N}{\log N}&\text{if }r=s=4\text{ and $\nu\notin 2\mathbb{Z}$}\\ \\ \asymp\dfrac{N}{(\log N)^{3/2}}&\text{if $\{r,s\}\neq\{4\}$ and $\{r,\nu\}\nsubseteq 2\mathbb{Z}$}\\ \\ \leq 1&\text{otherwise. }\end{cases}$
Case III. For $P_{3}$, from (2.5) we note that
$G=\gcd(2\mu(r-2),2\nu(s-2),\mu r,\nu s,\mu+\nu)=\gcd(4,r,s,\mu+\nu).$
Hence, $G=1$ if and only if either $r$ or $s$ of $\mu+\nu$ is odd. So, it follows from (2.7), Theorem A and Chebyshev’s theorem (Corollary 2.6 in [7]) that
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{2}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\hskip 5.0pt\begin{cases}\asymp\dfrac{N}{\log N}&\text{if }r=s=4\text{ and $\mu+\nu\notin 2\mathbb{Z}$}\\ \\ \asymp\dfrac{N}{(\log N)^{3/2}}&\text{if $\{r,s\}\neq\{4\}$ and $\{r,s,\mu+\nu\}\nsubseteq 2\mathbb{Z}$}\\ \\ \leq 1&\text{otherwise. }\end{cases}$
Since either $\nu$ or $\mu+\nu$ is odd for all mutually coprime pairs $\mu$ and $\nu$, the claim follows from the above cases. ∎

### v3

@@
 Abstract.
-For all integers $r,s>2$, there are infinitely many primes which can be written as the sum of an $r$-gonal number and an $s$-gonal number.
-In general, any linear combination of an $r$-gonal number and an $s$-gonal number with coprime positive integer coefficients produces infinitely many primes.
-However, unless $r=s=4$, such primes can not be described by finitely many congruences.
+For all integers $r,s>2$, each linear combination of an $r$-gonal number and an $s$-gonal number with coprime positive integer coefficients produces infinitely many primes.
+However, for any given pair of coprime positive integer coefficients of the polygonal numbers, the sum of the reciprocals of such primes diverges if and only if $r=s=4$.
 Key words and phrases:
@@
 1. Introduction and the main result
-A prime is a sum of two squares if and only if it is of the form $14k+$ 1 (see [8]).
+A prime is a sum of two squares if and only if it is of the form $14k+$ 1.
 Since there are infinitely many primes of the form $14k+$ 1, it follows that there are infinitely many primes which are sum of two squares.
-It is natural to wonder whether the same holds for triangular numbers or more generally, whether for all integers $r,s\geq 3$, there are infinitely many primes which can be represented as the sum of an $r$-gonal number and an $s$-gonal number, where the $m$-th $r$-gonal number is defined by (see Chapter 2 in [2])
+Since the sums of couples of triangular numbers are intimately related with the sums of couples of squares, viz.
+$n$ is a sum of two triangular numbers if and only if $14n+$ 1 is a sum of two squares, it is natural to expect that the primes which are sums of two triangular numbers will also have an equally appealing description in terms of some arithmetic progressions.
+However, it turns out not to be the case.
+In fact, the sum of the reciprocals of such primes converges.
+One may still wonder whether there are infinitely many such primes or more generally, whether for a given pair of integers $r,s>2$, there are infinitely many primes which are sum of an $r$-gonal number and an $s$-gonal number, where the $m$-th $r$-gonal number is defined by (see [2])
 $T(r,m):=(r-2)\frac{m(m-1)}{2}+m.$ (1.1)
-Even more generally, one may ask whether for each pair of integers $r,s\geq 3$ and for each pair of mutually coprime positive integers $\mu$ and $\nu$, there are infinitely many primes of the form
-$\mu T(r,m)+\nu T(s,n).$
-We answer this in affirmative.
+Even more generally, one may ask whether for each pair of integers $r,s>2$ and for every pair of mutually coprime positive integers $a$ and $b$, there are infinitely many primes of the form $aT(r,m)+bT(s,n).$ We answer this in the affirmative.
 Theorem 1.
-For sufficiently large $N$, for each pair of integers $r,s\geq 3$ and for each pair of mutually coprime positive integers $\mu$ and $\nu$, the number of primes of the form
-$\mu T(r,m)+\nu T(s,n)$
+Given a pair of integers $r,s>2$ and a pair of mutually coprime positive integers $a$ and $b$, for all $m,n\in\mathbb{N}$, the number of primes of the form
+$aT(r,m)+bT(s,n)$ (1.2)
 which are less than $N$ is
 $\asymp\begin{cases}\dfrac{N}{\log N},&\text{ if }r=s=4\\ \\ \dfrac{N}{(\log N)^{3/2}}&\text{otherwise.}\end{cases}$
-By an admissible arithmetic progression, we mean an integer arithmetic progression that has a pair of mutually coprime terms.
+From the above theorem, via Abel’s summation formula, we obtain
 Corollary 1.
-Unless $r=s=4$, for each pair of mutually coprime positive integers $\mu$ and $\nu$, the primes of the form
-$\mu T(r,m)+\nu T(s,n),$
-can not be described as the primes contained in a set of admissible arithmetic progressions.
-In other words, one can only describe the primes of the form
-$ax^{2}+by^{2}$
-by a finite set of congruences for certain $a,b\in\mathbb{N}$ (see [3]).
-However, Theorem 1 together with the Prime Number Theorem for arithmetic progressions (Corollary 11.20 in [7]) imply that for any pair of $a,b\in\mathbb{N}$, if we replace either $x^{2}$ or $y^{2}$ above by another polygonal number, the resulting linear combination of polygonal numbers misses infinitely many primes in every admissible arithmetic progression.
-In particular, the sequence $\{4k+1\}_{k\in\mathbb{N}}$ or its subprogressions are the only admissible arithmetic progressions such that each prime in those progressions is a sum of two polygonal numbers.
-It follows from Theorem 1 and Chebyshev’s theorem (Corollary 2.6 in [7]) that the set of the primes, which are sum of two polygonal numbers (or more generally, linear combination of two polygonal numbers with coprime positive integer coefficients) at least one of which is not a square, does not have positive relative upper density with respect to the set of all primes.
-So, Green-Tao Theorem [4] which guarantees the existence of arbitrarily long arithmetic progressions among the subsets of the primes with positive relative upper densities, is not applicable here either.
-From Theorem 1, for $\mu=\nu=1$ and $r=s=3$, we obtain the following corollary which answers the Open Problem 1 in [1].
+Given $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, the sum of the reciprocals of the primes of the form $aT(r,m)+bT(s,n)$ diverges if and only if $r=s=4$.
+In particular, Theorem 1 solves an open problem in [1], which asks how many primes are sums of two triangular numbers.
 Corollary 2.
 There are infinitely many primes which are sum of two triangular numbers.
-Since a number $n$ is a sum of two triangular numbers if and only if $14n+$ 1 is a sum of two squares (see [1]), we have
+In fact, there are infinitely many such primes in each of the residue classes ±$\pm 1$ modulo $4$.
+Theorem 2.
+Given two mutually coprime positive integers $a$ and $b$, the number of the primes in each of the residue classes ±$\pm 1$ modulo $4$, which are less than $N$ and which are linear combinations of two odd-sided polygonal numbers with coefficients $a$ and $b$ is
+$\asymp\dfrac{N}{(\log N)^{3/2}}.$
+Since a number $n$ is a sum of two triangular numbers if and only if $14n+$ 1 is a sum of two squares (see [1]) and since a prime is a sum of two squares if and only if it is of the form $14k+$ 1 (see [8]), we have
 Corollary 3.
-There are infinitely many primes $p$ such that $14p+$ 1 is a sum of two squares.
-It is natural to wonder whether there are infinitely many primes $p$ such that both $p$ and $14p+$ 1 are sums of two squares or more generally, whether both the residue classes $1$ and $3$ modulo $4$ contain infinitely many primes $p$ such that $14p+$ 1 is a sum of two squares.
-Even more generally, one may ask:
+The number of primes $p\leq N$ such that both $p$ and $14p+$ 1 are a sums of two squares is
+$\asymp\dfrac{N}{(\log N)^{3/2}}.$
+By an admissible arithmetic progression, we mean an integer arithmetic progression that has a pair of mutually coprime terms.
+Now, looking at Theorem 2, one may wonder whether every admissible arithmetic progression contains infinitely many primes which are sums of two triangular numbers.
+However, that does not hold in general!
+For example, if $n$ is congruent to $5$ or $8$ modulo 9, then $14n+$ 1 is not a sum of two squares.
+So, none of the arithmetic progressions $\{9k+5\}_{k\in\mathbb{N}}$ and $\{9k+8\}_{k\in\mathbb{N}}$ contains a sum of two triangular numbers.
+Further, Theorem 1 together with the Prime Number Theorem for arithmetic progressions (Corollary 11.20 in [7]) imply the following.
+Corollary 4.
+For each pair of integers $r,s>2$, at least one of which is not equal to $4$ and for every pair of $a,b\in\mathbb{N}$, the double sequence
+$\{aT(r,m)+bT(s,n)\}_{m,n\in\mathbb{N}}$
+misses infinitely many primes from every admissible arithmetic progression.
+In other words, one can only describe the primes of the form $am^{2}+bn^{2}$ by a finite set of congruences for certain $a,b\in\mathbb{N}$ (see [3]).
+However, for any pair of $a,b\in\mathbb{N}$, if we replace either $m^{2}$ or $n^{2}$ above by another polygonal number, the resulting linear combination of polygonal numbers fails to produce infinitely many primes contained in any given admissible arithmetic progression.
+Theorem 1 along with Chebyshev’s theorem (Corollary 2.6 in [7]) and Green-Tao Theorem [4] implies:
+Corollary 5.
+For each pair of coprime positive integers $a$ and $b$, there are arbitrariliy long arithmetic progressions among the primes of the form
+$am^{2}+bn^{2}.$
+It also follows from Theorem 1 and Chebyshev’s theorem that the sequence of the primes, which are linear combinations (with coprime positive integer coefficients) of two polygonal numbers at least one of which is not a square, has zero relative density with respect to the sequence of all primes.
+So, possible existence of arbitrary long arithmetic progressions among such primes (see the following table) is undetectable by Green-Tao Theorem [4].
+Table 1. Examples of arithmetic progressions among the primes which are sums of two triangular numbers
+Length Arithmetic progressions
+3 31, 37, 43
+4 43, 61, 79, 97
+5 61, 151, 241, 331, 421
+6 7, 37, 67, 97, 127, 157
+7 157, 37747, 75337, 112927, 150517, 188107, 225697
+8 3307, 69457, 135607, 201757, 267907, 334057, 400207, 466357
+9 3823, 6133, 8443, 10753, 13063, 15373, 17683, 19993, 22303
+10 73091, 135461, 197831, 260201, 322571, 384941, 447311, 509681, 572051, 634421
+The above table suggests the existence of arbitrarily long arithmetic progressions among the primes which are sums of two triangular number.
+More generally, we ask whether there are arbitrarily long arithmetic progressions among the primes which are linear combinations of two polygonal numbers with coprime positive integer coefficients.
 Open Question 1.
-Does every admissible arithmatic progression contain infinitely many primes which are sum of two triangular numbers?
-2. Proof of the main result
+For each choice of $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, are there are arbitrarily long arithmetic progressions among the primes of the form $aT(r,m)+bT(s,n)?$
+2. Proofs of the theorems
 The main ingredients of the proof of Theorem 1 are a couple of results by Iwaniec [5], which we state below.
 Theorem A (Theorem 1 in [5]).
-A Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g$ be of an irreducible polynomial of degree $2$ in $\mathbb{Z}[x,y]$ such that $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly independent and $P(x,y)$ represents arbitrarily large odd numbers for $x,y\in\mathbb{N}$.
+A Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g$ be of an irreducible polynomial in $\mathbb{Z}[x,y]$ such that $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly independent and $P(x,y)$ represents arbitrarily large odd numbers for $x,y\in\mathbb{N}$.
 Then
@@
 In particular, if the greatest common divisors of the coefficients of $P_{1}$, $P_{2}$, $P_{3}$ are 1, then each of these polynomials represents arbitrarily large odd numbers.
-In the following we compute the gcd $G$ of the coefficients of $P_{i}$ for $i\in\{1,2,3\}$.
+In the following we compute the gcd $G$ of the coefficients of $P_{1}$, $P_{2}$ and $P_{3}$.
+For $i\in\{1,2,3\}$ and $N\in\mathbb{N}$, let
+$\pi_{i}(N):=\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{i}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1.$
 Case I. For $P_{1}$, from (2.3) we note that
@@
 Hence, $G=1$ if and only if either $\mu r$ or $\nu s$ is odd. In particular, $G=1$ implies $D>0$ (see (2.7)). So, it follows from Theorem A that
-$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{1}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\hskip 5.0pt\begin{cases}\asymp\dfrac{N}{(\log N)^{3/2}}&\text{if $\{\mu r,\nu s\}\nsubseteq 2\mathbb{Z}$}\\ \\ \leq 1&\text{otherwise.
-}\end{cases}$
+$\pi_{1}(N)\asymp\begin{cases}\dfrac{N}{(\log N)^{3/2}}&\text{if $\{\mu r,\nu s\}\nsubseteq 2\mathbb{Z}$}\\ \\ 1&\text{otherwise. }\end{cases}$
 Case II. For $P_{2}$, from (2.4) we note that
@@
 Hence, $G=1$ if and only if either $r$ or $\nu$ is odd. So, it follows from (2.7), Theorem A and Chebyshev’s theorem (Corollary 2.6 in [7]) that
-$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{2}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\hskip 5.0pt\begin{cases}\asymp\dfrac{N}{\log N}&\text{if }r=s=4\text{ and $\nu\notin 2\mathbb{Z}$}\\ \\ \asymp\dfrac{N}{(\log N)^{3/2}}&\text{if $\{r,s\}\neq\{4\}$ and $\{r,\nu\}\nsubseteq 2\mathbb{Z}$}\\ \\ \leq 1&\text{otherwise.
+$\pi_{2}(N)\asymp\begin{cases}\dfrac{N}{\log N}&\text{if }r=s=4\text{ and $\nu\notin 2\mathbb{Z}$}\\ \\ \dfrac{N}{(\log N)^{3/2}}&\text{if $\{r,s\}\neq\{4\}$ and $\{r,\nu\}\nsubseteq 2\mathbb{Z}$}\\ \\ 1&\text{otherwise.
 }\end{cases}$
@@
 Hence, $G=1$ if and only if either $r$ or $s$ of $\mu+\nu$ is odd. So, it follows from (2.7), Theorem A and Chebyshev’s theorem (Corollary 2.6 in [7]) that
-$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{2}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\hskip 5.0pt\begin{cases}\asymp\dfrac{N}{\log N}&\text{if }r=s=4\text{ and $\mu+\nu\notin 2\mathbb{Z}$}\\ \\ \asymp\dfrac{N}{(\log N)^{3/2}}&\text{if $\{r,s\}\neq\{4\}$ and $\{r,s,\mu+\nu\}\nsubseteq 2\mathbb{Z}$}\\ \\ \leq 1&\text{otherwise.
+$\pi_{3}(N)\asymp\begin{cases}\dfrac{N}{\log N}&\text{if }r=s=4\text{ and $\mu+\nu\notin 2\mathbb{Z}$}\\ \\ \dfrac{N}{(\log N)^{3/2}}&\text{if $\{r,s\}\neq\{4\}$ and $\{r,s,\mu+\nu\}\nsubseteq 2\mathbb{Z}$}\\ \\ 1&\text{otherwise.
 }\end{cases}$
 Since either $\nu$ or $\mu+\nu$ is odd for all mutually coprime pairs $\mu$ and $\nu$, the claim follows from the above cases. ∎
+[…]
+Acknowledgments
+All the computations for Table 1 was done using PARI/GP.
+Also, Bhattacharya thankfully acknowledges the support from the Department of Science and Technology, Govt.
+of India through the INSPIRE Faculty Award IFA17-MA104 and Rahaman thankfully acknowledges the support from the Ministry of Education, Govt.
+of India through the Prime Minister Research Fellowship 0501972.

### v4

@@
 Abstract.
-For all integers $r,s>2$, each linear combination of an $r$-gonal number and an $s$-gonal number with coprime positive integer coefficients produces infinitely many primes.
-However, for any given pair of coprime positive integer coefficients of the polygonal numbers, the sum of the reciprocals of such primes diverges if and only if $r=s=4$.
+A linear combination $aT_{r}(m)+bT_{s}(n)$ of an $r$-gonal number $T_{r}(m)$ and an $s$-gonal number $T_{s}(n)$ with mutually coprime positive integer coefficients $a$ and $b$ produces infinitely many primes as $m$ and $n$ varies over the natural numbers, whereas the sum of the reciprocals of such primes converges unless $T_{r}(m)=m^{2}$ and $T_{s}(n)=n^{2}$.
+For each pair of coprime positive integers $a$ and $b$, there are arbitrary long arithmetic progressions among the primes of the form $am^{2}+bn^{2}$.
 Key words and phrases:
@@
 1. Introduction and the main result
-A prime is a sum of two squares if and only if it is of the form $14k+$ 1.
-Since there are infinitely many primes of the form $14k+$ 1, it follows that there are infinitely many primes which are sum of two squares.
+A prime is a sum of two squares if and only if it is of the form $4k+1$.
+Since there are infinitely many primes of the form $4k+1$, it follows that there are infinitely many primes which are sum of two squares.
 Since the sums of couples of triangular numbers are intimately related with the sums of couples of squares, viz.
-$n$ is a sum of two triangular numbers if and only if $14n+$ 1 is a sum of two squares, it is natural to expect that the primes which are sums of two triangular numbers will also have an equally appealing description in terms of some arithmetic progressions.
+$n$ is a sum of two triangular numbers if and only if $4n+1$ is a sum of two squares, it is natural to expect that the primes which are sums of two triangular numbers will also have an equally appealing description in terms of some arithmetic progressions.
 However, it turns out not to be the case.
@@
 One may still wonder whether there are infinitely many such primes or more generally, whether for a given pair of integers $r,s>2$, there are infinitely many primes which are sum of an $r$-gonal number and an $s$-gonal number, where the $m$-th $r$-gonal number is defined by (see [2])
-$T(r,m):=(r-2)\frac{m(m-1)}{2}+m.$ (1.1)
-Even more generally, one may ask whether for each pair of integers $r,s>2$ and for every pair of mutually coprime positive integers $a$ and $b$, there are infinitely many primes of the form $aT(r,m)+bT(s,n).$ We answer this in the affirmative.
+$T_{r}(m)\;:=\;(r-2)\frac{m(m-1)}{2}+m.$ (1.1)
+Even more generally, one may ask whether for each pair of integers $r,s>2$ and for every pair of mutually coprime positive integers $a$ and $b$, there are infinitely many primes of the form $aT_{r}(m)+bT_{s}(n).$ We answer this in the affirmative.
 Theorem 1.
 Given a pair of integers $r,s>2$ and a pair of mutually coprime positive integers $a$ and $b$, for all $m,n\in\mathbb{N}$, the number of primes of the form
-$aT(r,m)+bT(s,n)$ (1.2)
+$aT_{r}(m)+bT_{s}(n)$ (1.2)
 which are less than $N$ is
 $\asymp\begin{cases}\dfrac{N}{\log N},&\text{ if }r=s=4\\ \\ \dfrac{N}{(\log N)^{3/2}}&\text{otherwise.}\end{cases}$
+Recall that $f\asymp g$ if and only if there exists a pair of constants $C,C^{\prime}>0$ such that $Cg\leq f\leq C^{\prime}g$.
+In the first case above, we have $T_{4}(m)=m^{2}$.
+The surprise is in the other case, where even if one of the summands is a multiple of a square, the corresponding set of primes has a lower rate of growth.
 From the above theorem, via Abel’s summation formula, we obtain
 Corollary 1.
-Given $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, the sum of the reciprocals of the primes of the form $aT(r,m)+bT(s,n)$ diverges if and only if $r=s=4$.
-In particular, Theorem 1 solves an open problem in [1], which asks how many primes are sums of two triangular numbers.
+Given $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, the sum of the reciprocals of the primes of the form $aT_{r}(m)+bT_{s}(n)$ diverges if and only if $r=s=4$.
+In particular, Theorem 1 answers a question from [1], which asks how many primes are sums of two triangular numbers.
 Corollary 2.
 There are infinitely many primes which are sum of two triangular numbers.
-In fact, there are infinitely many such primes in each of the residue classes ±$\pm 1$ modulo $4$.
+In fact, there are infinitely many such primes in each of the residue classes $\pm 1$ modulo $4$.
 Theorem 2.
-Given two mutually coprime positive integers $a$ and $b$, the number of the primes in each of the residue classes ±$\pm 1$ modulo $4$, which are less than $N$ and which are linear combinations of two odd-sided polygonal numbers with coefficients $a$ and $b$ is
+Given two mutually coprime positive integers $a$ and $b$, the number of the primes in each of the residue classes $\pm 1$ modulo $4$, which are less than $N$ and which are linear combinations of two odd-sided polygonal numbers with coefficients $a$ and $b$ is
 $\asymp\dfrac{N}{(\log N)^{3/2}}.$
-Since a number $n$ is a sum of two triangular numbers if and only if $14n+$ 1 is a sum of two squares (see [1]) and since a prime is a sum of two squares if and only if it is of the form $14k+$ 1 (see [8]), we have
-Corollary 3.
-The number of primes $p\leq N$ such that both $p$ and $14p+$ 1 are a sums of two squares is
+In particular, for $a=b=1$ and for the primes congruent to $1$ modulo $4$, we have the following.
+Corollary.
+For each odd integer $r>1$, the number of primes $p\leq N$, which are sums of two $r$-gonal numbers and also sums of two squares, is
+$\asymp\frac{N}{(\log N)^{3/2}}.$
+Since a number $n$ is a sum of two triangular numbers if and only if $4n+1$ is a sum of two squares*** It follows from (1.1) that for all $r$, we have $4(r-2)(T_{r}(m)+T_{r}(n))+(r-4)^{2}\;=\;\left((r-2)(m-n)\right)^{2}+\left((r-2)(m+n)-r+4\right)^{2}.$ Conversely, if $4(r-2)M\;+\;(r-4)^{2}\;=\;A^{2}\;+\;B^{2}$ such that $A\;\pm\;B\;\equiv\;r\pmod{2(r-2)}$, then $M$ is a sum of two $r$-gonal numbers.
+(see [1]) and since a prime is a sum of two squares if and only if it is of the form $4k+1$ (see [8]), it follows that
+Corollary.
+The number of primes $p\leq N$ such that both $p$ and $4p+1$ are sums of two squares is
 $\asymp\dfrac{N}{(\log N)^{3/2}}.$
-By an admissible arithmetic progression, we mean an integer arithmetic progression that has a pair of mutually coprime terms.
-Now, looking at Theorem 2, one may wonder whether every admissible arithmetic progression contains infinitely many primes which are sums of two triangular numbers.
+By a primitive arithmetic progression, we mean an integer arithmetic progression that has a pair of mutually coprime terms.
+Now, looking at Theorem 2, one may wonder whether every primitive arithmetic progression contains infinitely many primes which are sums of two triangular numbers.
 However, that does not hold in general!
-For example, if $n$ is congruent to $5$ or $8$ modulo 9, then $14n+$ 1 is not a sum of two squares.
+For example, if $n$ is congruent to $5$ or $8$ modulo 9, then $4n+1$ is not a sum of two squares.
 So, none of the arithmetic progressions $\{9k+5\}_{k\in\mathbb{N}}$ and $\{9k+8\}_{k\in\mathbb{N}}$ contains a sum of two triangular numbers.
-Further, Theorem 1 together with the Prime Number Theorem for arithmetic progressions (Corollary 11.20 in [7]) imply the following.
+This prompts us to refine our query as follows:
+Question 1.
+Given $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, if there is no congruence obstruction, does every primitive arithmetic progression contain infinitely many primes which are of the form $aT_{r}(m)+bT_{s}(n)?$
+Theorem 1 together with the Prime Number Theorem for arithmetic progressions (Corollary 11.20 in [7]) imply the following.
 Corollary 3.
 For each pair of integers $r,s>2$, at least one of which is not equal to $4$ and for every pair of $a,b\in\mathbb{N}$, the double sequence
-$\{aT(r,m)+bT(s,n)\}_{m,n\in\mathbb{N}}$
-misses infinitely many primes from every admissible arithmetic progression.
+$\{aT_{r}(m)+bT_{s}(n)\}_{m,n\in\mathbb{N}}$
+misses infinitely many primes from every primitive arithmetic progression.
 In other words, one can only describe the primes of the form $am^{2}+bn^{2}$ by a finite set of congruences for certain $a,b\in\mathbb{N}$ (see [3]).
-However, for any pair of $a,b\in\mathbb{N}$, if we replace either $m^{2}$ or $n^{2}$ above by another polygonal number, the resulting linear combination of polygonal numbers fails to produce infinitely many primes contained in any given admissible arithmetic progression.
+However, for any pair of $a,b\in\mathbb{N}$, if we replace either $m^{2}$ or $n^{2}$ above by another polygonal number, the resulting linear combination of polygonal numbers fails to produce infinitely many primes contained in any given primitive arithmetic progression.
 Theorem 1 along with Chebyshev’s theorem (Corollary 2.6 in [7]) and Green-Tao Theorem [4] implies:
@@
 Table 1. Examples of arithmetic progressions among the primes which are sums of two triangular numbers
-Length Arithmetic progressions
+Length
+Arithmetic progressions
 3 31, 37, 43
@@
 9 3823, 6133, 8443, 10753, 13063, 15373, 17683, 19993, 22303
-10 73091, 135461, 197831, 260201, 322571, 384941, 447311, 509681, 572051, 634421
-The above table suggests the existence of arbitrarily long arithmetic progressions among the primes which are sums of two triangular number.
+10
+73091, 135461, 197831, 260201, 322571, 384941, 447311,
+509681, 572051, 634421
+The following table suggests the existence of arbitrarily long arithmetic progressions among the primes which are sums of two triangular number.
 More generally, we ask whether there are arbitrarily long arithmetic progressions among the primes which are linear combinations of two polygonal numbers with coprime positive integer coefficients.
-Open Question 1.
-For each choice of $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, are there are arbitrarily long arithmetic progressions among the primes of the form $aT(r,m)+bT(s,n)?$
+Question 2.
+For each choice of $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, are there are arbitrarily long arithmetic progressions among the primes of the form $aT_{r}(m)+bT_{s}(n)?$
 2. Proofs of the theorems
-The main ingredients of the proof of Theorem 1 are a couple of results by Iwaniec [5], which we state below.
+The main ingredients of the proof of Theorem 1 are a couple of results by Iwaniec [5], which we state below. A
 Theorem A (Theorem 1 in [5]).
-A Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g$ be of an irreducible polynomial in $\mathbb{Z}[x,y]$ such that $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly independent and $P(x,y)$ represents arbitrarily large odd numbers for $x,y\in\mathbb{N}$.
+Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g$ be an irreducible polynomial in $\mathbb{Z}[x,y]$ such that $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly independent and $P(x,y)$ represents arbitrarily large odd numbers for $x,y\in\mathbb{N}$.
 Then
-$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\gg\frac{N}{\log N},$ (2.1)
+$\Bigg{(}\sum_{\begin{subarray}{c}\text{prime\hskip 3.0pt}p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P(x,y)\text{\hskip 3.0ptfor $x,y\in\mathbb{N}$}\end{subarray}}1\Bigg{)}\;\;\gg\;\;\frac{N}{\log N},$ (2.1)
 if $\Delta:=b^{2}-4ac$ is a perfect square or $D:=af^{2}-bef+ce^{2}+g\Delta=0$, and
-$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\asymp\frac{N}{(\log N)^{3/2}},$ (2.2)
+$\Bigg{(}\sum_{\begin{subarray}{c}\text{prime\hskip 3.0pt}p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P(x,y)\text{\hskip 3.0ptfor $x,y\in\mathbb{N}$}\end{subarray}}1\Bigg{)}\;\;\asymp\;\;\frac{N}{(\log N)^{3/2}},$ (2.2)
 if $\Delta$ is not a perfect square and $D\neq 0$.
 Lemma B (Lemma 1 in [5]).
-B Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g\in\mathbb{Z}[x,y]$. Then $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly dependent if and only if
+B Let $P(x,y)\;=\;ax^{2}+bxy+cy^{2}+ex+fy+g$. Then $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly dependent if and only if
 $\Delta\;=\;b^{2}-4ac\;=\;0,\ \alpha\;:=\;bf-2ce\;=\;0\text{ \ and \ }\beta\;:=\;be-2af\;=\;0.$
-Proof of Theorem 1: For a pair of integers $r,s\geq 3$ and for a pair of mutually coprime integers $\mu,\nu>0$, considering without loss of generality all possible parities of the natural numbers $m$ and $n$ in
-$\mu T(r,m)+\nu T(s,n),$
+Proof of Theorem 1.
+For a pair of integers $r,s\geq 3$ and for a pair of mutually coprime integers $\mu,\nu>0$, considering without loss of generality all possible parities of the natural numbers $m$ and $n$ in
+$\mu T_{r}(m)+\nu T_{s}(n),$
 we note that (see (1.1)) the above linear combination is equal to
 $P_{1}(x,y)\;:=\;2\mu(r-2)x^{2}+2\nu(s-2)y^{2}+\mu(4-r)x+\nu(4-s)y$ (2.3)
-for $m=2x$ and $n=2y$,
+for m= 2xm\;=\;2x and n= 2yn\;=\;2y,
 $P_{2}(x,y)\;:=\;2\mu(r-2)x^{2}+2\nu(s-2)y^{2}+\mu(4-r)x+\nu sy+\nu$ (2.4)
-for $m=2x$ and $n=2y+1$,
+for m= 2xm\;=\;2x and n= 2y+1n\;=\;2y+1,
 $P_{3}(x,y)\;:=\;2\mu(r-2)x^{2}+2\nu(s-2)y^{2}+\mu rx+\nu sy+\mu+\nu$ (2.5)
-for $m=2x+1$ and $n=2y+1$.
+for m= 2x+1m\;=\;2x+1 and n= 2y+1n\;=\;2y+1.
 Since the coefficients of $x^{2}$ and $y^{2}$ in each of the above polynomials are positive and since the coefficient of $xy$ is zero, it follows that
@@
 Since $\mu,\nu>0$ and $r,s>2$, the first two equalities imply that $aa^{\prime}>0$ and $bb^{\prime}>0$ and the last equality implies that there exists $k\in\mathbb{Q}$ such that $\frac{a}{b}=-\frac{a^{\prime}}{b^{\prime}}=k$.
-Hence, substituting $kb$ for $a$ and $-kb^{\prime}$ for $a^{\prime}$ in the first equality and dividing both sides of it with the corresponding sides of the second equality above, we obtain
+Hence, substituting $kb$ for $a$ and $kb^{\prime}$ for $a^{\prime}$ in the first equality and dividing both sides of it with the corresponding sides of the second equality above, we obtain
 $k^{2}\;=\;-\frac{\mu(r-2)}{\nu(s-2)}<0,$
@@
 For $i\in\{1,2,3\}$ and $N\in\mathbb{N}$, let
-$\pi_{i}(N):=\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{i}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1.$
+$\pi_{i}(N)\;:=\;\sum_{{\begin{subarray}{c}\text{prime\hskip 3.0pt}p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{i}(x,y)\text{\hskip 3.0ptfor $x,y\in\mathbb{N}$}\end{subarray}}}1.$
 Case I. For $P_{1}$, from (2.3) we note that
@@
 Acknowledgments
+The authors would like to thank Prof.
+Don Zagier for his comments and suggestions.
 All the computations for Table 1 was done using PARI/GP.